OpenStudy (anonymous):

z = y/x if z = 4.0 and y = 6.8 what is the value of x ?

OpenStudy (irishboy123):

\(\large z = \frac{y}{x} \) ?? multiply both side by x and then divide both sides by z to get something that says \(x = ...\)

OpenStudy (anonymous):

by multuplying both sides by x still gives me 4.0 ?

OpenStudy (arindameducationusc):

just put the values 4=6.8/x =>x=4/6.8 Now You can simplify for your answer

OpenStudy (anonymous):

I got .588 , im rounding to the nearest tenth so .59. is this correct

OpenStudy (radar):

If you followed the instructions provided by IrishBoy123 You would of come up with a correct (and different) answer.

OpenStudy (radar):

If you followed those instructions, you would have isolated the x to the left-hand side of the equal sign and had a fraction consisting of y as the numerator and z the denominator. Try again.